What is ΔSsys for a freezing phase transition at 55.5 °C for a compound that freezes at 55.5 °C and the ΔHsys = -13.71 kJ mol-1 for this process?

Express your answer in J mol-1 K-1.
Answer #1

And :- At freezing point the change in entropy of system for freezing phase transistion is negative but entropy of surrounding increases as energy is released by system so overall entropy of universe increases and process is spontaneous at that temperature .

Given ∆Hsys= -13.71 KJmol-1

Freezing temperature ,T = 273+ 55 = 328K

Now , ∆S sys = ∆Hsys / T

∆Ssys = -13.71× 1000/(273+55) J mol-1 K-1

∆Ssys = -41.79 J mol-1 K-1.